Circular No. 6797

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 971227 L. Piro and P. Soffitta, Istituto Astrofisica Spaziale, CNR, Rome; R.

C. Butler, Agenzia Spaziale Italiana, Rome; L. A. Antonelli, F. Fiore, and M. Capalbi, BeppoSAX Science Data Center, Rome; A. Tesseri and V. Torroni, BeppoSAX Science Operation Center, Telespazio, Rome; and C. De Libero, BeppoSAX, Operation Control Center, Telespazio, Rome, report on behalf of the BeppoSAX team: “A follow-up observation of GRB 971227 was carried out with BeppoSAX 14 hr after the initial burst. Two previously unknown x-ray sources were detected within the WFC error circle (IAUC 6796) in the MECS x-ray telescopes. 1SAX J1257.5+5915 is located at α = 12^{h}57^{m}33^{s}, δ = +59^{o}15′27′′ (equinox 2000.0; error radius ∼ 1′.5, including systematics in the position reconstruction and statistical errors). Three radio sources of the NRAO/VLA Sky Survey catalogue (cf.

http://www.cv.nrao.edu/~jcondon/nvss.html) lie within this error cir-

cle; the strongest one has a flux of 20 mJy at 1.4 GHz and is polarized. The x-ray source does not show any significant variation during the 20-hr observation. The average countrate in the range 1.3–10 keV was 0.0012 ± 0.0004 counts/s for two MECS units, corresponding to a 2–10-keV flux of 1 × 10^{−13} erg cm^{−2} s^{−1}. The other source, 1SAX J1257.3+5924, was detected at α = 12^{h}57^{m}15^{s}, δ = +59^{o}24′02′′. The countrate in the first 8 hr of the observation was 0.0032 ± 0.001 counts/s, which corresponds to a 2–10-keV flux of 3 × 10^{−13} erg cm^{−2} s^{−1}. In the following part of the observation, the source was not detected, with a 3σ upper limit of 0.002 counts/s. We conclude that 1SAX J1257.3+5924 is the best candidate for the x-ray afterglow of GRB 971227.”

SUPERNOVA 1997ef IN UGC 4107 J. Y. Wei, J. Y. Hu, Y. L. Qiu, and Q. Y. Qiao, Beijing Astronomical

Observatory; and J. S. Deng and J. H. You, Shanghai Jiao Tong University, report the following R magnitudes for SN 1997ef: Dec. 4.8 UT, 16.4; 11.7, 16.1; 12.7, 16.1; 13.8, 16.2; 16.8, 16.3; 22.8, 16.3; 24.7, 16.2; 26.8, 16.2.

COMET C/1995 O1 (HALE-BOPP) Visual m_{1} estimates: Nov. 2.74 UT, 5.9 (T. Karhula, Yerecoin, W. Aus-

tralia, naked eye); 18.52, 6.6 (D. Seargent, The Entrance, N.S.W., 10×50 binoculars); Dec. 4.25, 7.0 (J. de S. Aguiar, Campinas, Brazil, 11×80 binoculars); 29.53, 7.6 (A. Pearce, Nedlands, W. Australia, 20×80 binoculars).

1997 December 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green